Two objects with masses of m_1 = 2.00 kg and m_2 =
Answer #1

Given masses,

m1 = 2kg

m2 =8 kg

First, write a force balance on the 2kg object:

The forces are gravity Fg (downward) and tension T (upward), and the object is moving from the problem statement:

m*a = T - mg , or

(2kg)*a = T - (2kg)(9.8m/s^2)

Do the same for the 8 kg object:

m*a = T - mg, or

(8kg)*a = T - (8kg)*(9.8m/s^2)

Since the two blocks are connected via a taut string, a is the same (in magnitude) for both but opposite in sign. We can subtract the two equations to eliminate T from the equations and solve for a:

(2kg + 8kg)*a = (-2kg + 8kg)*(9.8m/s^2)

a= (-2kg + 8kg)*(9.8m/s^2)/(10kg) = 5.886 m/s^2 (PART-B ANSWER) (upward for 2kg block and downward for 8 kg block)

You can substitute this back into the first equation to solve for T:

T = (2kg)*(5.886m/s^2) + (2kg)(9.8m/s^2) = 31.392 N

TENSION IN THE STRING = 31.392 N (PART-A ANSWER)

For part c) you can use the equation of motion:

d = 1/2*a*t^2

= (0.5)*(5.886m/s^2)*(1 s)^2 = 2.943 m (PART-C ANSWER) (upward for 2kg block, downward for 8kg block)